OpenStudy (anonymous):

the difference between two positive numbers is 5 and their product is 84. find the numbers

OpenStudy (anonymous):

b = a + 5 ab = 84 Now substitute "a+5" for "b" in the second equation and solve for a. Then plug that into the first equation to get b.

OpenStudy (saifoo.khan):

x - y = 5 xy = 84

OpenStudy (anonymous):

ah, thank you guys.

OpenStudy (anonymous):

\[\sqrt{84}=9.165 \]A guess that the larger is 3 more than 9, 12, and the lesser is, 5 less than 12 gives 7. Lucked out. 12*7 = 84